OpenStudy (anonymous):

help . . .where is the vertex of y=1/2x^2+2x-3?

OpenStudy (anonymous):

derivate this expression one time y'=2/2x+2 the vertex is for y'=0 so x=-2

OpenStudy (anonymous):

magic formula for vertex is \[-\frac{b}{2a}\]

OpenStudy (anonymous):

now solve y=1/2 x^2+2x-3 with x=-2 y=-5

OpenStudy (anonymous):

in this case \[a=\frac{1}{2},b=2\] so \[-\frac{b}{2a}=-\frac{2}{2\times \frac{1}{2}}=-2\]

OpenStudy (anonymous):

first coordinate of vertex is -2, second is what you get when you replace x by -2 in the original expression \[\frac{1}{2}\times (-2)^2+2\times -2-3\] \[2-4-3\] \[-5\] so vertex is \[(-2,-5)\]

OpenStudy (anonymous):

i suppose if you know calculus you can use it to find the vertex, but you certainly don't need it. just remember \[-\frac{b}{2a}\]

OpenStudy (anonymous):

thanks satellite!